OpenStudy (marcelie):

help please !!!! At a family reunion, there were only blood relatives, consisting of children, parents, and grandparents, in attendance. There were 400 people total. There were twice as many parents as grandparents, and 50 more children than parents. How many children, parents, and grandparents were in attendance?

Directrix (directrix):

Let G represent the number of Grandparents. There were twice as many parents as grandparents: Let 2G represent the number of Parents 50 more children than parents. Let 2G + 50 represent the Children ---------------------------- G + P + C = 400 Solve this equation for G: G + 2G + 2G + 50 + 400 @marcelie Post what you get, okay?

Directrix (directrix):

This above G + 2G + 2G + 50 + 400 should be G + 2G + 2G + 50 = 400 G + 2G + 2G + 50 + 400 5G + 50 = 400 5*G = 350

Directrix (directrix):

@marcelie What is 350 divided by 5 equal to?

OpenStudy (koikkara):

\(350/5=70\) as \(7(5)=35\)

OpenStudy (koikkara):

@Directrix

OpenStudy (marcelie):

sorry for the late response i got it . ty :)

Directrix (directrix):

You are welcome.